OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. 2 is a factor of n^2 - n + 2

OpenStudy (anonymous):

so i know i need to prove that s1 is correct as the first step

OpenStudy (anonymous):

1^2 - 1 + 2 1-1+2 0+2 2

OpenStudy (anonymous):

where do i go from here?

Nnesha (nnesha):

correct. now make an assumption assume its true for n=k and then we have to prove it's true for n=k+1 based on the assumption

OpenStudy (anonymous):

so (k+1)^2 - (k+1) + 2

Nnesha (nnesha):

yes first we assumed n =k so \[\rm k^2-k+2\] 2 is a factor of k^2-k+2 and then assumed n=k+1 so yes that's correct simplify that

OpenStudy (anonymous):

k^2-k+2 is already simplified

Nnesha (nnesha):

no i mean simplify this `(k+1)^2 - (k+1) + 2`

OpenStudy (anonymous):

k^2+1^2-k+1+2 k^2+1-k+3 k^2-k+4

Nnesha (nnesha):

ohhhh wait no! that's not how we should foil (k+1)^2\[\rm (k+1)^2 \cancel{=} k^2+1^2\]

Nnesha (nnesha):

if u need shortcut then take `square of first term + multiply both term by 2 + square of last term`

OpenStudy (anonymous):

lol, i forgot

OpenStudy (anonymous):

k^2+2k+1

Nnesha (nnesha):

or draw a box like this